OpenStudy (anonymous):

Solve the following equation: 1/x+2/(x-1)=3

OpenStudy (jamesj):

He doesn't need the answer; he needs detailed steps

OpenStudy (anonymous):

Thanks james

OpenStudy (anonymous):

then myininaya will write it out

myininaya (myininaya):

first of all we do not want x=0 or x=1 anyways ... multiply both sides by x(x-1) to clear the bottoms \[x(x-1)\frac{1}{x}+x(x-1)\frac{2}{x-1}=x(x-1)3\] \[(x-1)+2x=3x(x-1)\]

myininaya (myininaya):

we will use the distributive property on the right hand side \[(x-1)+2x=3x^2-3x\] combine like terms on left hand side \[3x-1=3x^2-3x\] put everything on one side

myininaya (myininaya):

\[3x^2-3x-3x+1=0\]

myininaya (myininaya):

\[3x^2-6x+1=0\]

OpenStudy (anonymous):

Thank you

myininaya (myininaya):

np